Sweetie145:

Solve for w.4(w − 12) = 16w =

Joe348:

do you mean 4(w-12)=16w?

Sweetie145:

@joe348 wrote:
do you mean 4(w-12)=16w?
yes

MusicGeek:

First you have to distribute w to what's in the parenthesis. So it would be 4w-48=16w. Then subtract 4w to both sides. -48=12w. Then you divide both sides by 12. Once you do that, you will have w=-4.

Joe348:

damn stealing my questions brah

MusicGeek:

Sorry dude. I just like helpin peoples as fast as i can

Sweetie145:

@musicgeek wrote:
First you have to distribute w to what's in the parenthesis. So it would be 4w-48=16w. Then subtract 4w to both sides. -48=12w. Then you divide both sides by 12. Once you do that, you will have w=-4.
thx so much🥰